Fortes Inc. has provided the following data concerning one of the products in its standard cost system. Variable manufacturing overhead is applied to products on the basis of direct labor-hours.

Inputs Standard Quantity or Hours per Unit of Output Standard Price or Rate
Direct materials 8.4 ounces $ 6.50 per ounce
Direct labor 0.6 hours $ 23.80 per hour
Variable manufacturing overhead 0.6 hours $ 4.80 per hour

The company has reported the following actual results for the product for April:

Actual output 6,000 units
Raw materials purchased 54,760 ounces
Actual cost of raw materials purchased $ 299,360
Raw materials used in production 50,410 ounces
Actual direct labor-hours 3,270 hours
Actual direct labor cost $ 80,810
Actual variable overhead cost $ 14,953

Required:

a. Compute the materials price variance for April.

b. Compute the materials quantity variance for April.

c. Compute the labor rate variance for April.

d. Compute the labor efficiency variance for April.

e. Compute the variable overhead rate variance for April.

f. Compute the variable overhead efficiency variance for April.

(Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)
Answer #1

a) Material price variance = (6.5*54760-299360) = 56580 F

b) Material quantity variance = (6000*8.4-50410)*6.5 = 65 U

C) Labor rate variance = (23.80*3270-80810) = 2984 U

d) Labor efficiency variance = (6000*.6-3270)*23.80 = 7854 F

e) Variable overhead rate variance = (4.8*3270-14953) = 743 F

f) Variable overhead efficiency variance = (3600-3270)*4.8 = 1584 F